Give the nuclear symbol (isotope symbol) for the isotope of bromine that contains 44 neutrons per atom.

nuclear symbol:

Give the nuclear symbol (isotope symbol) for the isotope of potassium that contains 20 neutrons per atom.

nuclear symbol:

Give the nuclear symbol (isotope symbol) for the isotope of silver that contains 60 neutrons per atom.

nuclear symbol:
Answer #1

Nuclear symbol is represented as,

where,

E = symbol for the element (H - Hydrogen, He - Helium, Br - Bromine, etc)

Z = Atomic number (For hydrogen H = ,1. Helium He = 2, bromine Br = 35)

M = Mass number. This is the sum of = no of protons + no of neutrons. This is the general convention of representing nuclear symbols.

The way to solve this questions is

1. Given the name of element, refer a periodic table and note its atomic number. The atomic number = number of protons = no of electrons (as an atom is electrically neutral). This is Z in the nuclear symbol.

2. Next, add the number of neutrons to the above atomic number. This will give the mass number M.

Thus for bromine, Z = 35 and given number of neutrons is 44. Then mass number M = 44+35=79. Thus nuclear symbol is

For Potassium, Z = 19, given no of neutrons 20, thus M = 19+20 = 39

For Silver, Z = 47, and given number of neutrons = 60, thus mass number M = 47+60 = 107